OpenStudy (supercait10):

Find the missing Length of the side of a square with an area of 121. Answers: A.484 B. 30.25 C. 11 D. 6.5 plz just give me the answer Medal and fan as a reward

OpenStudy (campbell_st):

a square has all sides equal. and if the length is l the area is \[A = l^2\] since you know the area \[121 = l^2\] just take the square root to side the side length

OpenStudy (anonymous):

just divide 121 by 4 and you get an answer of B

OpenStudy (supercait10):

484? and thanks for the help

OpenStudy (campbell_st):

its not 484

OpenStudy (campbell_st):

and its not 30.25 take the square root of 121

OpenStudy (supercait10):

11?

OpenStudy (campbell_st):

that's correct

OpenStudy (supercait10):

thanks for the help I might need more for others if thats okay

OpenStudy (anonymous):

hold on, can you explain to me how 11 is correct? a square has 4 equal sides, with an area of 121.

OpenStudy (campbell_st):

the question asks for the missing side length of a square that has an area of 121 sq units you answer 30.25 is correct if the perimeter is 121 \[A = 11 x 11 ~~~A = 121\]

OpenStudy (campbell_st):

if you use 30.25 as the side length the area is \[A = 30.25 \times 30.25 ~~~A = 915.0625\]

OpenStudy (er.mohd.amir):

area of square = side *side 121=side*side 11*11=side * side side=11

OpenStudy (anonymous):

okay yeah, i now see what mistake i made

OpenStudy (supercait10):

Okay if its no bother can you help me out on some more questions I have to get a test done before 6:30am this morning

OpenStudy (anonymous):

ill try helping now that i know where my mistake was

OpenStudy (anonymous):

it may take me a while tho because i am feeding by baby bunnies

OpenStudy (supercait10):

okay thats fine

OpenStudy (supercait10):

Identify all the sets to which the number belongs. Choose from Rational Number, Irrational number, whole number, and or a integer. 0.123333 A. Rational number B. Irrational number C.Integer, and rational number D. Whole number, Integer, and rational number

OpenStudy (anonymous):

i would say rational number

OpenStudy (supercait10):

okay

OpenStudy (anonymous):

okay